# Functors on the category of finite sets revisited

### v2

@@ 2.2. Projectives
The functor $P^{\mathbf{FA}}_{\mathbf{1}}$ identifies as $X\mapsto\mathbbm{k}[X]$, where $\mathbbm{k}[X]$ is the $\mathbbm{k}$-vector space freely generated by $X$; this comes with the canonical basis $\{[x]\ |\ x\in X\}$. Hence $P^{\mathbf{FA}}_{\mathbf{n}}$ identifies as the functor $X\mapsto\mathbbm{k}[X]^{\otimes n}$.
@@
This Proposition exhibits a left augmentation of the ‘unit’ $\mathbbm{k}{\mathbf{FI}}\hookrightarrow\mathbbm{k}\mathbf{FA}$ (corresponding to the canonical inclusion), in the terminology of Positselski [Pos21]. Left (respectively right) augmentations are exploited in [Pow24b], using results of Positselski.
@@
There is a ‘categorically dual’ version of Proposition 2.8 for surjective maps:
@@ 2.3. A norm-like map
and this is a morphism of $\mathbbm{k}\mathrm{Aut}(X)$-bimodules. This has adjoint
@@
which is a morphism of $\mathbbm{k}\mathrm{Aut}(X)^{\mathrm{op}}\otimes\mathbbm{k}\mathbf{FA}$-modules (for the action of $\mathbbm{k}\mathrm{Aut}(X)^{\mathrm{op}}$ on the codomain arising from the right regular action on $\mathbbm{k}\mathrm{Aut}(X)$).
@@
induced by the $\mathbbm{k}$-linear surjection $\mathbbm{k}\mathrm{Aut}(X)\twoheadrightarrow\mathbbm{k}$ that sends a generator $[\alpha]$ ($\alpha\in\mathrm{Aut}(X)$) to zero unless $\alpha=\mathrm{Id}_{X}$, when the image is $1$.
@@
For the final statement, using ${\mathbf{FS}}(X,X)$ as the canonical basis of $\mathbbm{k}{\mathbf{FS}}(X,X)$, the dual basis provides the isomorphism of $\mathbbm{k}$-vector spaces $D\mathbbm{k}{\mathbf{FS}}(X,X)\cong\mathbbm{k}{\mathbf{FS}}(X,X)$, so that the map identifies as a $\mathbbm{k}$-linear map $\mathbbm{k}\mathrm{Aut}(X)\rightarrow\mathbbm{k}\mathrm{Aut}(X)$. It is straightforward to check that this is given by $[\alpha]\mapsto[\alpha^{-1}]$, where $\alpha\in\mathrm{Aut}(X)$. In particular, this is an isomorphism, as required. ∎
@@
3. Analysing the standard projectives
The purpose of this section is to commence the analysis of the structure of the projectives of $\mathcal{F}(\mathbf{FA})$, working over an arbitrary field $\mathbbm{k}$. The key input is Proposition 3.9, which yields a more efficient set of projective generators of $\mathcal{F}(\mathbf{FA})$ (see Corollary 3.11).
3.1. Filtering the standard projectives
The projective $P^{\mathbf{FA}}_{\mathbf{0}}$ identifies as the constant functor with value $\mathbbm{k}$. This is not simple; there is a non-split short exact sequence
@@ 3.1. Filtering the standard projectives
where $\mathbbm{k}_{\mathbf{0}}$ is the functor supported on $\mathbf{0}$ with value $\mathbbm{k}$ and $\overline{\mathbbm{k}}\subset\mathbbm{k}$ is the subfunctor supported on non-empty finite subsets. It is easily checked that both $\mathbbm{k}_{\mathbf{0}}$ and $\overline{\mathbbm{k}}$ are simple $\mathbbm{k}\mathbf{FA}$-modules.
Now consider $P^{\mathbf{FA}}_{\mathbf{1}}$, the functor $X\mapsto\mathbbm{k}[X]$ ($\mathbbm{k}[X]$ is also sometimes written here as $\mathbbm{k}X$). By Yoneda, $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{1}},P^{\mathbf{FA}}_{\mathbf{0}})\cong\mathbbm{k}$, with generator
@@
(3.3) $0\rightarrow\overline{P}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\overline{\mathbbm{k}}\rightarrow 0$
@@
Were this short exact sequence to split, there would be a non-trivial surjection $P^{\mathbf{FA}}_{\mathbf{1}}\twoheadrightarrow\overline{P}$. However, $\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}_{\mathbf{1}},\overline{P})=0$, by Yoneda, since $\overline{P}(\mathbf{1})=0$ by Lemma 3.2. ∎
@@
One has the following alternative description of these tensor products:
@@
$0\rightarrow\overline{P}^{\otimes n}\rightarrow P^{\mathbf{FA}}_{\mathbf{n}}\rightarrow\bigoplus_{i=1}^{n}P^{\mathbf{FA}}_{\mathbf{n}\backslash\{i\}}$
where the right hand map is induced by the inclusions $\mathbf{n}\backslash\{i\}\subset\mathbf{n}$.
Proof.
For $n=1$, this is simply the definition of $\overline{P}$. The general result follows by considering the $n$-fold tensor product $(P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow P^{\mathbf{FA}}_{\mathbf{0}})^{\otimes n}$ of complexes. ∎
@@
3.2. First (non)splitting results
Lemma 3.8.
This yields the splitting result:
@@ 3.2. First (non)splitting results
If $P$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n}}$ for some $n>0$, then $P\otimes\overline{P}$ is projective in $\mathcal{F}(\mathbf{FA})$ and is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n+1}}$.
@@
Applying $P\otimes-$ to the short exact sequence (3.3) yields the short exact sequence
$0\rightarrow P\otimes\overline{P}\rightarrow P\otimes P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow P\otimes\overline{\mathbbm{k}}\rightarrow 0$
The hypothesis upon $P$ implies that $P(\mathbf{0})=0$ and hence that $P\otimes\overline{\mathbbm{k}}$ is isomorphic to $P$. Hence the short exact sequence in the statement splits, since $P$ is projective. The final statement follows by using the isomorphism $P^{\mathbf{FA}}_{\mathbf{n+1}}\cong P^{\mathbf{FA}}_{\mathbf{n}}\otimes P^{\mathbf{FA}}_{\mathbf{1}}$, so that $P\otimes P^{\mathbf{FA}}_{\mathbf{1}}$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n+1}}$ and is projective. ∎
Corollary 3.10.
For any $n\in\mathbb{N}$, the subfunctor $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$ of $P^{\mathbf{FA}}_{\mathbf{1}}\otimes(P^{\mathbf{FA}}_{\mathbf{1}})^{\otimes n}\cong P^{\mathbf{FA}}_{\mathbf{n+1}}$ is a direct summand and is projective. Hence there is an isomorphism:
$P^{\mathbf{FA}}_{\mathbf{n+1}}\cong P^{\mathbf{FA}}_{\mathbf{1}}\otimes\big{(}\overline{P}\oplus\overline{\mathbbm{k}}\big{)}^{\otimes n}\cong P^{\mathbf{FA}}_{\mathbf{1}}\otimes\bigoplus_{k=0}^{n}(\overline{P}^{\otimes k})^{\binom{n}{k}}.$
Proof.
The projectivity of $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$ and the splitting of $P^{\mathbf{FA}}_{\mathbf{n+1}}$ follow by induction from Proposition 3.9. ∎
This allows Proposition 3.6 to be refined:
Corollary 3.11.
The set of objects $\{P^{\mathbf{FA}}_{\mathbf{0}}=\mathbbm{k}\}\amalg\{P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\ |\ n\in\mathbb{N}\}$ is a set of projective generators of $\mathcal{F}(\mathbf{FA})$.
@@
By construction, there is a short exact sequence:
(3.4) $0\rightarrow\overline{P}^{\otimes n+1}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\rightarrow\overline{P}^{\otimes n}\rightarrow 0$
This does not split in general (if $\mathbbm{k}$ has characteristic zero, this follows from the results of Section 5), whence $\overline{P}^{\otimes n}$ is not projective.
@@
(3.5) $\ldots\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n+1}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\rightarrow\ldots\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}},$
where $n$ corresponds to the homological degree. Splicing with $0\rightarrow\overline{\mathbbm{k}}\rightarrow\mathbbm{k}=P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\mathbbm{k}_{\mathbf{0}}\rightarrow 0$ gives a projective resolution of $\mathbbm{k}_{\mathbf{0}}$.
The following gives further information on the surjection of Proposition 2.8:
@@
It suffices to show that, for $f:\mathbf{n}\hookrightarrow X$ an injection, the generator $[f]\in\mathbbm{k}{\mathbf{FI}}(\mathbf{n},X)$ lies in the image. Write $x_{i}:=f(i)$, for $i\in\mathbf{n}$. Then one has the element:
$[x_{1}]\otimes([x_{2}]-[x_{1}])\otimes\ldots\otimes([x_{n}]-[x_{1}])\in\big{(}P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n-1}\big{)}(X).$
@@
4. Further results on projectives in arbitrary characteristic
This section is a continuation of Section 3. Theorem 4.4 provides initial information on morphisms between functors of the form $\overline{P}^{\otimes n}$, proved using the idempotent constructed in Section 4.1. This is used to show that the indecomposable projectives in $\mathbbm{k}\mathbf{FA}$-modules have finite length (see Corollary 4.7).
Theorem 4.8 gives an interpretation of the right augmentation $\mathbbm{k}\mathbf{FA}\rightarrow\mathbbm{k}{\mathbf{FS}}$ within the theory, in terms of morphisms from the functors $\overline{P}^{\otimes n}$ to the standard projectives.
4.1. An explicit idempotent
The splitting given by Corollary 3.10 yields the idempotent $P^{\mathbf{FA}}_{\mathbf{n+1}}\twoheadrightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\hookrightarrow P^{\mathbf{FA}}_{\mathbf{n+1}}$. By Yoneda, this corresponds to an idempotent in the algebra $\mathbbm{k}\mathbf{FA}(\mathbf{n+1},\mathbf{n+1})$. An explicit idempotent is constructed below, refining the argument used in the proof of Proposition 3.13.
@@ 4.1. An explicit idempotent
    For a subset $Y\subset(\mathbf{n+1})\backslash\{1\}$, define $f_{Y}:\mathbf{n+1}\rightarrow\mathbf{n+1}$ to be the set map
    $i\mapsto\left\{\begin{array}[]{ll}1&i\in Y\\ i&i\not\in Y.\end{array}\right.$
@@
    For $n\in\mathbb{N}$, define $\pi_{\mathbf{n}}\in\mathbbm{k}\mathbf{FA}(\mathbf{n+1},\mathbf{n+1})$ by
    $\pi_{\mathbf{n}}:=\sum_{Y\subset(\mathbf{n+1})\backslash\{1\}}(-1)^{|Y|}[f_{Y}].$
Precomposing with $\pi_{\mathbf{n}}$ induces an endomorphism $P^{\mathbf{FA}}_{\mathbf{n+1}}\rightarrow P^{\mathbf{FA}}_{\mathbf{n+1}}$. Recall that $P^{\mathbf{FA}}_{\mathbf{n+1}}(X)$ identifies with $\mathbbm{k}[X]^{\otimes n+1}$, for a finite set $X$. This has basis indexed by $X^{n+1}$; a tuple $(x_{i})$ corresponds to the basis element $[x_{1}]\otimes\ldots\otimes[x_{n+1}]$. The action of $\pi_{\mathbf{n}}$ is given by
$[x_{1}]\otimes[x_{2}]\otimes\ldots\otimes[x_{n+1}]\mapsto[x_{1}]\otimes([x_{2}]-[x_{1}])\otimes\ldots\otimes([x_{n+1}]-[x_{1}]),$
where the right hand side lies in $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\subset P^{\mathbf{FA}}_{\mathbf{n+1}}$.
@@
For $n\in\mathbb{N}$, $\pi_{\mathbf{n}}$ is an idempotent in $\mathbbm{k}\mathbf{FA}(\mathbf{n+1},\mathbf{n+1})$ and induces the projection onto $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}\subset P^{\mathbf{FA}}_{\mathbf{n+1}}$.
@@
From the identification of the action of $\pi_{\mathbf{n}}$ preceding the statement, it is clear that $\pi_{\mathbf{n}}$ induces a surjection $P^{\mathbf{FA}}_{\mathbf{n+1}}\twoheadrightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$. It remains to show that $\pi_{\mathbf{n}}$ is an idempotent.
The key input is the obvious fact that, for $Y,Z\subset(\mathbf{n+1})\backslash\{\mathbf{1}\}$, one has $f_{Y}\circ f_{Z}=f_{Y\cup Z}=f_{Z}\circ f_{Y}$. This implies that
$\pi_{\mathbf{n}}\circ\pi_{\mathbf{n}}=\sum_{W\subset(\mathbf{n+1})\backslash\{\mathbf{1}\}}\Big{(}\sum_{\begin{subarray}{c}Y,Z\subset(\mathbf{n+1})\backslash\{\mathbf{1}\}\\ Y\cup Z=W\end{subarray}}(-1)^{|Y|+|Z]}[f_{W}]\Big{)}.$
Thus, to prove the result, we require to prove that $\sum_{\begin{subarray}{c}Y,Z\subset(\mathbf{n+1})\backslash\{\mathbf{1}\}\\ Y\cup Z=W\end{subarray}}(-1)^{|Y|+|Z]}=(-1)^{|W|}$. It is sufficient to show that, for a finite set $W$,
$\sum_{\begin{subarray}{c}Y,Z\subset W\\ Y\cup Z=W\end{subarray}}(-1)^{|Y|+|Z]}=(-1)^{|W|}.$
This is proved by increasing induction upon $|W|$. The case $W=\emptyset$ is immediate, serving as the initial case. For the inductive step, suppose the result holds for $W$ and consider $W_{+}=W\amalg\{w\}$. Using the notation $(-)_{+}$ and the indexing set of pairs $(Y,Z)$ for $W$, there are two mutually exclusive cases:
  1. (1)
    pairs $(Y_{+},Z_{+})$ (i.e., $w$ lies in both of the subsets);
  2. (2)
    pairs $(Y_{+},Z)$ or $(Y,Z_{+})$ (i.e., $w$ lies in only one of the subsets).
Using the inductive hypothesis and taking into account the signs, the first terms contribute $(-1)^{|W|}$; the second terms contribute $(-2)*(-1)^{|W|}$. Hence their sum gives $-(-1)^{|W|}=(-1)^{|W_{+}|}$, as required. ∎
Proposition 4.2 has the immediate consequence:

### v3

@@ 2.2. Projectives
The functor $P^{\mathbf{FA}}_{\mathbf{1}}$ identifies as $X\mapsto\mathbbm{k}[X]$, where $\mathbbm{k}[X]$ is the $\mathbbm{k}$-module freely generated by $X$; this comes with the canonical basis $\{[x]\ |\ x\in X\}$. Hence $P^{\mathbf{FA}}_{\mathbf{n}}$ identifies as the functor $X\mapsto\mathbbm{k}[X]^{\otimes n}$.
@@
This Proposition exhibits a left augmentation of the ‘unit’ $\mathbbm{k}{\mathbf{FI}}\hookrightarrow\mathbbm{k}\mathbf{FA}$ (corresponding to the canonical inclusion), in the terminology of Positselski [4]. Left (respectively right) augmentations are exploited in [6], using results of Positselski.
@@
There is a categorically dual version of Proposition 2.8 for surjective maps:
@@ 2.3. A norm-like map
To simplify the exposition, in this subsection we suppose that $\mathbbm{k}$ is a field. Vector space duality induces an exact functor $D:\mathcal{F}(\mathcal{C})^{\mathrm{op}}\rightarrow\mathcal{F}(\mathcal{C}^{\mathrm{op}})$ where, for a $\mathbbm{k}\mathcal{C}$-module $F$, the dual $DF$ is defined by $DF(X):=\mathrm{Hom}_{\mathtt{Mod}_{\mathbbm{k}}}(F(X),\mathbbm{k})$.
@@
and this is a morphism of $\mathbbm{k}\mathrm{Aut}(X)$-bimodules. By adjunction, this corresponds to a morphism
@@
of $\mathbbm{k}\mathrm{Aut}(X)^{\mathrm{op}}\otimes\mathbbm{k}\mathbf{FA}$-modules (for the action of $\mathbbm{k}\mathrm{Aut}(X)^{\mathrm{op}}$ on the codomain arising from the right regular action on $\mathbbm{k}\mathrm{Aut}(X)$).
@@
induced by the $\mathbbm{k}$-linear surjection $\mathbbm{k}\mathrm{Aut}(X)\twoheadrightarrow\mathbbm{k}$ that sends a generator $\alpha]$ ($\alpha\in\mathrm{Aut}(X)$) to zero unless $\alpha=\mathrm{Id}_{X}$, when the image is $1$.
@@
For the final statement, using ${\mathbf{FS}}(X,X)$ as the canonical basis of $\mathbbm{k}{\mathbf{FS}}(X,X)$, the dual basis provides the isomorphism of $\mathbbm{k}$-vector spaces $D\mathbbm{k}{\mathbf{FS}}(X,X)\cong\mathbbm{k}{\mathbf{FS}}(X,X)$, so that the map identifies as a $\mathbbm{k}$-linear map $\mathbbm{k}\mathrm{Aut}(X)\rightarrow\mathbbm{k}\mathrm{Aut}(X)$. It is straightforward to check that this is given by $\alpha]\mapsto[\alpha^{-1}]$, where $\alpha\in\mathrm{Aut}(X)$. In particular, this is an isomorphism, as required. ∎
@@
3. Filtering the standard projectives
In this section we commence the analysis of the structure of the projectives of $\mathcal{F}(\mathbf{FA})$, working over a unital commutative ring $\mathbbm{k}$. We first introduce the subfunctor $\overline{P}\subset P^{\mathbf{FA}}_{\mathbf{1}}$, whence the tensor products $\overline{P}^{\otimes t}$ and $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}$ for $t\in\mathbb{N}$. These are key players in the paper; in particular, Proposition 3.10 gives a simple direct proof of the fact that $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}$ is projective. In Section 4 it is explained how this yields a more efficient set of projective generators of $\mathcal{F}(\mathbf{FA})$.
The projective $P^{\mathbf{FA}}_{\mathbf{0}}$ identifies as the constant functor with value $\mathbbm{k}$. There is a non-split short exact sequence
@@ 3. Filtering the standard projectives
where $\mathbbm{k}_{\mathbf{0}}$ is the functor supported on $\mathbf{0}$ with value $\mathbbm{k}$ and $\overline{\mathbbm{k}}\subset\mathbbm{k}$ is the subfunctor supported on non-empty finite subsets.
Remark 3.1.
If $\mathbbm{k}$ is a field, then both $\mathbbm{k}_{\mathbf{0}}$ and $\overline{\mathbbm{k}}$ are simple $\mathbbm{k}\mathbf{FA}$-modules.
Now consider $P^{\mathbf{FA}}_{\mathbf{1}}$, the functor $X\mapsto\mathbbm{k}[X]$ ($\mathbbm{k}[X]$ may also be written as $\mathbbm{k}X$). By Yoneda, $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{1}},P^{\mathbf{FA}}_{\mathbf{0}})\cong\mathbbm{k}$, with generator
@@
(3.3) $0\rightarrow\overline{P}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\overline{\mathbbm{k}}\rightarrow 0.$
@@
Were the short exact sequence to split, there would be a non-trivial surjection $P^{\mathbf{FA}}_{\mathbf{1}}\twoheadrightarrow\overline{P}$. However, $\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}_{\mathbf{1}},\overline{P})=0$, by Yoneda, since $\overline{P}(\mathbf{1})=0$ by Lemma 3.3. ∎
@@
Lemma 3.6.
Then, for $n$ a positive integer, forming the complex $(P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\overline{\mathbbm{k}})^{\otimes n}$ gives the following generalization of the short exact sequence (3.3):
@@
$0\rightarrow\overline{P}^{\otimes n}\rightarrow P^{\mathbf{FA}}_{\mathbf{n}}\rightarrow\bigoplus_{i=1}^{n}P^{\mathbf{FA}}_{\mathbf{n}\backslash\{i\}}\rightarrow\bigoplus_{\begin{subarray}{c}T\subset\mathbf{n}\\ |T|=2\end{subarray}}P^{\mathbf{FA}}_{\mathbf{n}\backslash T}\rightarrow\ldots\rightarrow\overline{\mathbbm{k}}\rightarrow 0$
[…]
@@
Lemma 3.6 also yields the following projectivity result:
@@
If $Q$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n}}$ for some $n>0$, then $Q\otimes\overline{P}$ is projective in $\mathcal{F}(\mathbf{FA})$ and is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n+1}}$.
Hence, for any $t\in\mathbb{N}$, the functor $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}$ is projective.
@@
Applying $Q\otimes$- to the short exact sequence (3.3) yields the short exact sequence
$0\rightarrow Q\otimes\overline{P}\rightarrow Q\otimes P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow Q\otimes\overline{\mathbbm{k}}\rightarrow 0.$
The hypothesis upon $Q$ implies that $Q(\mathbf{0})=0$ and hence that $Q\otimes\overline{\mathbbm{k}}$ is isomorphic to $Q$. Hence the short exact sequence in the statement splits, since $Q$ is projective. The final statement follows by using the isomorphism $P^{\mathbf{FA}}_{\mathbf{n+1}}\cong P^{\mathbf{FA}}_{\mathbf{n}}\otimes P^{\mathbf{FA}}_{\mathbf{1}}$, so that $Q\otimes P^{\mathbf{FA}}_{\mathbf{1}}$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n+1}}$ and is projective.
The projectivity of $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}$ then follows by an obvious induction upon $t$. ∎
@@
An alternative proof of the projectivity of $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}$ is given in Section 4, using the identification given in Proposition 4.14.
Remark 3.12.
By construction (see Corollary 4.16), for $t\in\mathbb{N}$, there is a short exact sequence:
(3.4) $0\rightarrow\overline{P}^{\otimes t+1}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}\rightarrow\overline{P}^{\otimes t}\rightarrow 0.$
This does not split in general (if $\mathbbm{k}$ has characteristic zero, this follows from the results of Section 5), whence $\overline{P}^{\otimes t}$ is not projective.
@@
(3.5) $\ldots\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t+1}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes t}\rightarrow\ldots\rightarrow P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}\rightarrow P^{\mathbf{FA}}_{\mathbf{1}},$
where $t$ corresponds to the homological degree. Splicing with $0\rightarrow\overline{\mathbbm{k}}\rightarrow\mathbbm{k}=P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\mathbbm{k}_{\mathbf{0}}\rightarrow 0$ gives a projective resolution of $\mathbbm{k}_{\mathbf{0}}$.
Truncating by omitting terms in homological degree less than $n$, for $n$ a positive integer, yields a projective resolution of $\overline{P}^{\otimes n}$.
For later use, we record the following, which gives further information on the surjection of Proposition 2.8:
@@
It suffices to show that, for $f:\mathbf{n}\hookrightarrow X$ an injection, the generator $f]\in\mathbbm{k}{\mathbf{FI}}(\mathbf{n},X)$ lies in the image. Write $x_{i}:=f(i)$, for $i\in\mathbf{n}$. Then one has the element:
$x_{1}]\otimes([x_{2}]-[x_{1}])\otimes\ldots\otimes([x_{n}]-[x_{1}])\in\big{(}P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n-1}\big{)}(X).$
@@
4. Splitting the projectives and applications
The purpose of this section is to revisit the projectives of the form $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes n}$ and establish some of their properties. For this, in Section 4.1, the projective direct summands $\mathbf{Q}^{X,x}_{\emptyset}$ of $P^{\mathbf{FA}}_{X}$ are introduced, where $x\in X$. Their significance is established by Corollary 4.8 which shows how they yield a set of projective generators of $\mathcal{F}(\mathbf{FA})$. A counterpart is established for $\mathcal{F}(\mathbf{FA}^{\mathrm{op}})$.
Section 4.2 makes the structure of these projectives more explicit, in particular establishing the relation with the projectives introduced in Section 3. This analysis is essential input to the proof of Theorem 4.19 in Section 4.3.
These results are then used in Section 4.4 to study morphisms from $\overline{P}^{\otimes n}$ to the standard projectives $P^{\mathbf{FA}}_{\mathbf{t}}$, for $t\in\mathbb{N}$, together with its applications. The main result of this subsection is Theorem 4.23.
Finally, in Section 4.5 these results are applied to establish some basic finiteness properties, also giving some initial information on the simple $\mathbbm{k}\mathbf{FA}$-modules.
Throughout this section, unless indicated otherwise, $\mathbbm{k}$ is a unital, commutative ring.
4.1. Splitting $P^{\mathbf{FA}}_{X}$ using Möbius inversion
In this section a splitting of $P^{\mathbf{FA}}_{X}$ is given by using the general, conceptual framework of Möbius inversion.
Remark.
This approach (together with indications of proofs) was suggested by the referee, together with ingredients feeding into Theorem 4.19, which was proposed by the referee.
We work with finite pointed sets of the form $(X,x)$, so that $X$ is a finite set and $x\in X$. (This will allow the naturality of the constructions to be analysed should that be required.)
@@ 4.1. Splitting $P^{\mathbf{FA}}_{X}$ using Möbius inversion
Suppose that $(X,x)$ is a pointed finite set and $T,Z\subseteq X\backslash\{x\}$ are subsets.
@@
[…]
    $s^{X,x}_{T}\colon y\mapsto\left\{\begin{array}[]{ll}x&y\in T\\ y&\mbox{otherwise.}\end{array}\right$.
@@
    Define $f^{X,x}_{T}\in\mathbf{FA}(X,X)$ by $f^{X,x}_{T}:=i^{X}_{T}\circ s^{X,x}_{T}$.
  3. (3)
    Define $e^{X,x}_{Z}\in\mathbbm{k}\mathbf{FA}(X,X)$ by
    $e^{X,x}_{Z}:=\sum_{Z\subseteq Y\subseteq X\backslash\{x\}}(-1)^{|Y\backslash Z|}[f^{X,x}_{Y}].$
The following records some basic properties of these morphisms for later use.
Lemma 4.2.
For $(X,x)$ a finite pointed set and subsets $T,T^{\prime}\subseteq X\backslash\{x\}$,
  1. (1)
[…]
  2. (2)
    $f^{X,x}_{T}\circ f^{X,x}_{T^{\prime}}=f^{X,x}_{T\cup T^{\prime}}=f^{X,x}_{T^{\prime}}\circ f^{X,x}_{T}$ in $\mathbf{FA}(X,X)$;
  3. (3)
    $f^{X,x}_{T}\circ e^{X,x}_{T}=e^{X,x}_{T}$ in $\mathbbm{k}\mathbf{FA}(X,X)$;
  4. (4)
[…]
Proof.
Point (1) is clear and (2) is a straightforward verification. The property (3) is established readily by using (2) and property (4) is again a straightforward verification. ∎
The significance of the $e^{X,x}_{T}$ is established by the following, which is an immediate consequence of the Möbius inversion result [1, Theorem A.1].
@@
For $(X,x)$ a finite pointed set, $e^{X,x}_{T}\mid\emptyset\subseteq T\subseteq X\backslash\{x\}\}$ is a set of orthogonal idempotents that sum to $\mathrm{Id}_{X}]\in\mathbbm{k}\mathbf{FA}(X,X)$.
Remark 4.4.
[…]
Definition 4.5.
For $(X,x)$ a finite pointed set and $T\subseteq X\backslash\{x\}$, define $\mathbf{Q}^{X,x}_{T}$ by
$\mathbf{Q}^{X,x}_{T}:=\mathbbm{k}\mathbf{FA}(X,\bullet)e^{X,x}_{T}\in\mathrm{Ob}\hskip 2.0pt\mathcal{F}(\mathbf{FA}).$
By Proposition 4.3, there is a direct sum decomposition as a finite direct sum of projectives in $\mathcal{F}(\mathbf{FA})$:
(4.1) $P^{\mathbf{FA}}_{X}\cong\bigoplus_{\emptyset\subseteq Y\subseteq X\backslash\{x\}}\mathbf{Q}^{X,x}_{Y}.$
One can reduce to considering the projectives of the form $\mathbf{Q}^{X\backslash Y,x}_{\emptyset}$ using the following morphisms.
Definition 4.6.
For $(X,x)$ a pointed finite set and $T\subseteq X\backslash\{x\}$, let
  1. (1)
[…]
  2. (2)
[…]
These morphisms relate the idempotents $e^{X,x}_{T}$ and $e^{X\backslash T,x}_{\emptyset}$ as follows:
Proposition 4.7.
For $(X,x)$ a pointed finite set and $T\subseteq X\backslash\{x\}$, there are equalities:
[…]
$a^{X,x}_{T}\circ b^{X,x}_{T}$ $=$ $e^{X,x}_{T}\mbox{ in }\mathbbm{k}\mathbf{FA}(X,X).$
These induce an isomorphism $\mathbf{Q}^{X,x}_{T}\cong\mathbf{Q}^{X\backslash T,x}_{\emptyset}$.
@@
Using Lemma 4.2, we have equalities
$b^{X,x}_{T}\circ a^{X,x}_{T}$ $=$ $e^{X\backslash T,x}_{\emptyset}\circ s^{X,x}_{T}\circ e^{X,x}_{T}\circ i^{X}_{T}\circ e^{X\backslash T,x}_{\emptyset}$
$=$ $e^{X\backslash T,x}_{\emptyset}\circ s^{X,x}_{T}\circ i^{X}_{T}\circ e^{X\backslash T,x}_{\emptyset}\circ e^{X\backslash T,x}_{\emptyset}$
$=$ $e^{X\backslash T,x}_{\emptyset}.$
Likewise, we have equalities
$a^{X,x}_{T}\circ b^{X,x}_{T}$ $=$ $e^{X,x}_{T}\circ i^{X}_{T}\circ e^{X\backslash T,x}_{\emptyset}\circ s^{X,x}_{T}\circ e^{X,x}_{T}$
$=$ $e^{X,x}_{T}\circ e^{X,x}_{T}\circ i^{X}_{T}\circ s^{X,x}_{T}\circ e^{X,x}_{T}$
$=$ $e^{X,x}_{T}\circ e^{X,x}_{T}\circ f^{X,x}_{T}\circ e^{X,x}_{T}.$
Since $f^{X,x}_{T}\circ e^{X,x}_{T}=e^{X,x}_{T}$, by Lemma 4.2, the second equality follows.
Thus these induce isomorphisms
${\mathbf{Q}^{X\backslash T,x}_{\emptyset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a^{X,x}_{T}}\scriptstyle{\cong}$${\mathbf{Q}^{X,x}_{T}.\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{b^{X,x}_{T}}$